OpenStudy (anonymous):

from integers 1 to 30,how many ways are there to add 3 integers that is divisible by 3? help!!

OpenStudy (sriram):

are we allowed to add one integer twice

OpenStudy (anonymous):

yup.

jhonyy9 (jhonyy9):

1. add to 3 one 3 = 6 the first nr again 3 = 9 - the second and again 3 = 12 the 3th number 2. multiply 3 with 3 = 9 this 9 with 2 = 18 the second and to this assum 3=21 the 3th

jhonyy9 (jhonyy9):

but you can us the exponential possibillity 3*1 =3 3*2 =9 3*3 =27

jhonyy9 (jhonyy9):

or you can us the numbers multiply of 9 like 9 and 18 and 27

jhonyy9 (jhonyy9):

or check thos numbers like 12 --- 1+2=3 --- divisidily with 3 15 --- 1+5=6 --- divisibily with 3 18 --- 1+8=9 --- div with 3 21 24 27

OpenStudy (sriram):

Every natural number can be represented as 3n 3n+1 or 3n+2 here n is any number(1,2,3......) in our question we need n from 0 to 10 (to get 1to 30) for any number to be divisible by 3 it has to be of the form 3n+3n or 3n+1 + 3n+2 Extending this concept to three numbers it has to be of the form 3n+3n+3n or 3n+ 3n+1 +3n +2 taking the first case there are 10 numbers of the form 3n (3,6,9......30) as repeatition is allowed it can be got by 10 10 10 C + C + C =10 +10 + 10 =30 1 1 1 Second case 3n +(3n+1) +(3n+2) 3n---10 numbers of this type (3,6,9...30) 3n+1-----9 numbers of this type (1,4,7....28) 3n+2 ------9 numbers of this type (2,5,8,....29) this can be got by 10 9 9 C + C + C = 10 + 9 + 9 =28 1 1 1 therefore total ways possible =30+28=58 jst check wheather it is correct coz i could have made some mistakes. sorry couldn't explain it better on comp.

OpenStudy (anonymous):

thanks SRIRAM!!

OpenStudy (sriram):

welcome